Everton college paid $1525 for an order of 45 calculators. The store paid $10 for each scientific calculator. The others, all others cost $53 each. How many of each type was ordered? __________Scientific __________graphing.
Answer by jorel1380(3719)

m+n=45
10m+53n=1525
10m+10n=450
43n=1075
n=25
m=20
Everton bought 20 scientific calculators and 25 graphing calculators..